# Avat3r: Large Animatable Gaussian Reconstruction Model for High-fidelity 3D Head Avatars

### v1

@@ 4.5 Ablations
In the following, we study the efficacy of our design choices for Avat3r. In particular, we look at the usefulness of DUSt3R’s position maps, the Sapiens feature maps, and using input images with different expressions during training. Our findings are summarized in Tab. 2 and Fig. 6.
@@
A model that is trained on only multi-view consistent input images produces slightly sharper images. However, any inconsistencies in the input images, that can easily happen in practice, carry over to the generated 3D representation causing unwanted artifacts.
@@
Figure 7: Application scenarios of Avat3r. We show-case various use-cases of our model such as animating a randomly sampled 3D head, animating an image generated by a text-to-image diffusion model, and animating a picture of a greek bust.
@@ 4.6 Applications
Finally, we show some applications of Avat3r in Fig. 7. We study three use-cases: (i) Generating a novel 3D head with GGHead [27] and then animating it with Avat3r, (ii) Generating a portrait image with a text-to-image diffusion model [9], lifting it with a 3D GAN [55], and animating it with Avat3r, and (iii) Lifting an existing photograph with a 3D GAN [55] and animating it with Avat3r. The results demonstrate that, despite being designed for a few-shot scenario and only being trained on real studio data, Avat3r can be used in a variety of casual situations. For example, our findings for case (i) suggest that Avat3r may be used to animate images from any 3D GAN that generates 3D heads, effectively making any purely static 3D generation model animatable.
@@ 4.7 Limitations
Despite Avat3r’s great ability to create 3D head avatars from a few or even just a single input image, several limitations remain. For example, the current pipeline for inferring avatars from a single image requires a 3D GAN for 3D lifting which accumulates errors in the process in two ways: First, the 3D GAN won’t be able to perfectly explain the person, inevitably loosing some details, and second, existing inversion frameworks employ NeRF-based 3D GANs that rely on screen-space super-resolution after rendering which introduces view-inconsistencies to the final images. Both aspects hamper the quality of the final avatar. However, these issues will decrease as newer and better 3D GANs become available that do not require screen-space rendering anymore [27, 44]. Another limitation of our framework is the requirement of camera poses during inference. While reasonable estimates for face images can be obtained by head tracking algorithms [59], incorrect camera estimates still remain a potential source of error for real use-cases. Possible remedies may include teaching the network to deal with incorrect camera estimates already during training, or dropping the requirement for camera poses alltogether [45]. Finally, our current pipeline does not provide control over the lighting and instead bakes in light effects from the input images. This limits applications that put avatars into novel environments since the generated head will look out of place. Adressing this issue by explicitly disentangling the lighting properties is an exciting avenue for future research.
@@ Acknowledgements
Appendix A Analysis of Data Efficiency
In Fig. 8 we show how our model scales with the number of training subjects and input views. For the latter, we disable DUSt3R as it produces less reliable position maps for 2 input views, and cannot be executed at all for 1 input view. We see a clear improvement when using more training subjects as well as using more input views. However, further scaling the number of input views also has drawbacks, as it drastically increases runtime due to dense attention inside the transformer and the increased number of Gaussians that have to be rendered. We qualitatively analyze the effect of using more train subjects in Fig. 9 and the effect of the number of input views in Fig. 10.
@@ Appendix A Analysis of Data Efficiency
Appendix B Effect of Skip Connections
@@
Appendix C Training Details
@@
Head-centric coordinates.
k-farthest viewpoint sampling.
Input timestep sampling.
Speed.
Appendix D Hyperparameters

### v2

@@ 4.4 Ablations
In the following, we study the efficacy of our design choices for Avat3r. In particular, we look at the usefulness of DUSt3R’s position maps, the Sapiens feature maps, and using input images with different expressions during training. More ablations and a runtime analysis can be found in the supplemental material. Our findings are summarized in Tab. 2 and Fig. 4.
@@
A model that is trained on only multi-view consistent input images produces slightly sharper images. However, any inconsistencies in the input images, which can easily happen in practice, carry over to the generated 3D representation causing unwanted artifacts.
4.5 Additional Results
Extreme expressions.
We showcase the generalization capability of our method by interpolating between two extreme expressions in Fig. 5. Both source and driver person are unseen.
Driver 1 $\overrightarrow{\hskip 108.405pt\text{\footnotesize{Expression Interpolation}}\hskip 108.405pt}$ Driver 2
Figure 5: Expression Generalization.
Adapting to arbitrary input views.
Despite only being trained with 4 input views, our model can be directly used when more or fewer views are available, see Fig. 6.
[…]
input images 3 4 (train) 5 6
Figure 6: Input Flexibility. Our model accepts different numbers of views. Colors indicate from which view a Gaussian originates.
@@ 4.5 Additional Results
Figure 7: Application scenarios of Avat3r. We demonstrate various use-cases of our model such as animating a randomly sampled 3D head, animating an image generated by a text-to-image diffusion model, and animating a picture of a greek bust.
@@ 4.6 Applications
To make Avat3r amenable for inference on only a single input image, we make use of a pre-trained 3D GAN [77] to first lift the single image to 3D and then render four views of the head. In Fig. 7, we show three use-cases of Avat3r: (i) Generating a novel 3D head with GGHead [37] and then animating it with Avat3r, (ii) Generating a portrait image with a text-to-image diffusion model [15], lifting it with a 3D GAN [77], and animating it with Avat3r, and (iii) Lifting an existing photograph with a 3D GAN [77] and animating it with Avat3r. The results demonstrate that, despite being designed for a few-shot scenario and only being trained on real studio data, Avat3r can be used in a variety of casual situations. For example, our findings for case (i) suggest that Avat3r may be used to animate images from any 3D GAN that generates 3D heads, effectively making any purely static 3D generation model animatable.
@@ 4.7 Limitations
Despite Avat3r’s great ability to create 3D head avatars from a few or even just a single input image, several limitations remain. For example, the current pipeline for inferring avatars from a single image requires a 3D GAN for 3D lifting which accumulates errors in the process in two ways: First, the 3D GAN cannot perfectly explain the person, inevitably loosing some details, and second, existing inversion frameworks employ NeRF-based 3D GANs that rely on screen-space super-resolution after rendering which introduces view-inconsistencies to the final images. Both aspects hamper the quality of the final avatar. However, these issues will decrease as newer and better 3D GANs become available that do not require screen-space rendering anymore [37, 63]. Another limitation of our framework is the requirement of camera poses during inference. While reasonable estimates for face images can be obtained by head tracking algorithms [84], incorrect camera estimates still remain a potential source of error for real use-cases. Possible remedies may include teaching the network to deal with incorrect camera estimates already during training, or dropping the requirement for camera poses alltogether [64]. Finally, our current pipeline does not provide control over the lighting and instead bakes in light effects from the input images. This limits applications that put avatars into novel environments since the generated head will look out of place. Adressing this issue by explicitly disentangling the lighting properties is an exciting avenue for future research.
@@ Acknowledgements
[…]
Ava256 1 GAGAvatar 18.1 0.66 0.37 7.0 0.45
1 GAGAvatar 19.6 0.68 0.31 5.6 0.33
1 Ours3DGAN 19.1 0.68 0.39 6.0 0.37
1 Ours1 19.6 0.70 0.46 5.7 0.40
NeRSemble Video FlashAvatar 15.0 0.42 0.61 8.8 0.17
4 Ours 20.5 0.75 0.33 3.7 0.50
4 Ours+ 984 21.6 0.77 0.29 3.8 0.76
1 HeadNeRF 9.7 0.69 0.48 5.2 0.18
1 Portrait4D-v2 17.5 0.58 0.36 5.4 0.41
1 GAGAvatar 18.7 0.70 0.35 5.4 0.44
1 GAGAvatar 18.9 0.70 0.32 4.8 0.24
1 Ours3DGAN 19.5 0.71 0.38 4.3 0.31
1 Ours3DGAN + 984 19.4 0.71 0.38 4.6 0.46
1 Ours1 19.8 0.73 0.38 4.4 0.30
re-trained on Ava256 1trained on 1 input view
Table 3: Quantitative Comparison on 3D head avatar creation from various. Our approach performs competitively when only a single input image is available, despite not being designed for a single image use-case. It also performs much better compared to monocular methods that receive a full video as input. Ours3DGAN denotes a 4-shot model where the 4 required images are obtained from the single input via 3D lifting with a 3D GAN. Ours+ 984 denotes a model that is additionally fine-tuned on 984 neutral identities for better identity preservation (CSIM metric).
Appendix A Additional Results
A.1 Avat3rs from Phone Captures and Accessories
Fig. 12 showcases additional phone scans captured by users on their own devices, including challenging cases with glasses, rotated inputs, and a NeRSemble subject wearing a headscarf. Despite not being trained on accessories, the model handles glasses reasonably well and reconstructs the headscarf with high fidelity.
FlashAvatar frontal FlashAvatar Ours GT
Figure 9: Comparison with FlashAvatar on NeRSemble.
Input GAGavatar GAGavatar Ours1 Ours3DGAN GT
Figure 10: Single-image comparison on Ava256. We compare Avat3r with the recent 3D-aware portrait animation method GAGAvatar [13] in a self-reenactment scenario on hold-out persons from the Ava256 dataset. GAGavatar denotes a version of the baseline that we trained on the Ava256 dataset. Ours1 is a version of our model that was trained on only 1 input image (see Appendix B). Our method with 3D lifting (Ours3DGAN) shows better rendering quality than the baselines, especially for extreme expressions and viewing angles.
Input HeadNeRF Portrait4Dv2 GAGavatar GAGavatar Ours1 Ours3DGAN GT
Figure 11: Single-image comparison on NeRSemble. We compare Avat3r with the recent 3D-aware portrait animation methods GAGAvatar [13] (GAGavatar denotes a version of the baseline that we trained on the Ava256 dataset) and Portrait4Dv2 [18] as well as the NeRF-based face model HeadNeRF [27] in a self-reenactment scenario on persons from the NeRSemble dataset [35] Note that the NeRSemble dataset has not been used during training and therefore constitutes an evaluation scenario where both source and driver image are out-of-domain. Ours1 is a version of our model that was trained on only 1 input image (see Appendix B). Our method with 3D lifting (Ours3DGAN) shows better rendering quality than the baselines in these challenging scenarios where input and target view are from opposite sides of the face.
Figure 12: More Avat3r reconstruction results from phone scans with different phone cameras and persons with challenging clothing.
A.2 Single-shot 3D Head Avatar Creation
To make Avat3r amenable for inference on only a single input image, we make use of a pre-trained 3D GAN [77] to first lift the single image to 3D and then render four views of the head. These renderings then constitute the input for Avat3r. We conduct comparisons with the recent 3D-aware portrait animation method GAGAvatar [13]. Specifically, we compare with two version of GAGAvatar: One provided by the authors which is trained on VFHQ [68], and another version, denoted as GAGAvatar, that we train on the Ava256 dataset in the same setting as our method. To drive GAGAvatar, we use their monocular FLAME tracker to obtain tracked meshes. We also compare with another 3D portrait animation method, Portrait4Dv2 [18], and HeadNeRF [27]. Fig. 10 and Fig. 11 show qualitative comparisons between our method and the baselines for single input images of hold out persons. Note that our method performs competitively compared to the single-input baselines despite never being trained for a single-shot scenario. We also include a version of our model that was trained on a single input image without DUSt3R. In general, we find that Avat3r produces more realistic facial expressions than GAGAvatar and Portrait4Dv2 which are limited by FLAME’s expression space. Furthermore, our method allows much more extreme viewpoint changes without sacrificing rendering quality. Tab. 3 shows quantitative results. Note that, in contrast to portrait animation methods like GAGAvatar and Portrait4Dv2, our method can benefit when more input views are available (see Appendix B and tab. 1 in the main paper).
A.3 Comparison with Monocular Methods
We compare with the recent monocular approach FlashAvatar and provide it with a full video sequence from NeRSemble. As shown in Fig. 9, the monocular method performs well from the training view but fails on novel views due to overfitting. Please also refer to the Gaussian Avatar Fusion or HeadGAP papers, which made similar observations. In contrast, our learned reconstruction prior produces a plausible 3D head avatar. Quantitative comparisons are shown in Tab. 3.
@@ A.3 Comparison with Monocular Methods
Appendix B Analysis of Data Efficiency
B.1 Scaling Subjects and Views from Ava256
In Fig. 13 we show how our model scales with the number of training subjects and input views available in the Ava256 dataset. For the analysis on the number of input views, we disable DUSt3R as it produces less reliable position maps for 2 input views, and cannot be executed at all for 1 input view. We see a clear improvement when using more training subjects as well as using more input views. However, further scaling the number of input views also has drawbacks, as it drastically increases runtime due to dense attention inside the transformer and the increased number of Gaussians that have to be rendered. We qualitatively analyze the effect of using more train subjects in Fig. 14 and the effect of the number of input views in Fig. 15.
B.2 Effect of Adding More Neutral Subjects
Ours3DGAN Ours3DGAN + 984 Ours3DGAN Ours3DGAN + 984
Figure 16: Improved identity preservation by including 984 additional identities with neutral expression during training.
Since Avat3r is only trained on 244 subjects from the Ava256 dataset, it is at risk of overfitting to those identities during training time. In our experiments, we observe that while there is a slight identity shift between the final avatar and the person in the input images, the expression transfer works quite well. We therefore hypothesize that for further improvements on the model’s generalization capabilities it is not necessary to also add thousands of expressions for each additional person. To test this, we fine-tune our model with 984 additional identities from an internal dataset — just one expression each — adding only 0.08% to the training data. As shown in Fig. 16 and confirmed by improved CSIM metrics in Tab. 3, this small addition noticeably improves identity retention for these challenging cases.
[…]
in [s] in [fps]
HeadNeRF 65 111
Portrait4D-v2 0.2 4
GAGAVatar 0.1 63
GPAvatar 0.2 9.5
Ours1 1.1 53
Ours3DGAN 17.9 7.9
Ours 12.3 7.9
Table 4: Runtime analysis. Our method can create a high-quality avatar in a few seconds, and animate it at interactive rates.
Appendix C Inference efficiency & driving speed.
While Avatar creation takes several seconds for our method due to DuSt3R and Sapiens, we can cache all activations up to the final cross-attention layers afterwards, leading to expression driving at 7.9fps for our 4-shot model and 53fps for our single-shot model, see Tab. 4. Runtimes measured on a single RTX3090 GPU.
@@ B.2 Effect of Adding More Neutral Subjects
Input Images Animated 3D Head Avatars
Appendix D Effect of Skip Connections
Appendix E Training Details
@@
Head-centric coordinates.
Expression codes.
Our architecture is agnostic to the specific choice of animation signal. For experiments on ava256, we used the dataset’s expression codes that were originally predicted by a generalized expression encoder providing a driving signal beyond FLAME’s topology. For experiments on NeRSemble and in-the-wild driving videos, we fine-tuned our model using FLAME codes obtained by running GAGAvatar’s version of Metrical Tracker [84, 13]. This shows that Avat3r learns a general notion of facial expressions that can be adapted to fit a specific driving signal.
k-farthest viewpoint sampling.
Input timestep sampling.
Speed.
Appendix F Hyperparameters